Activity 4

The Math Club organized a fund-raising activity for the support of the students who cannot avail cellphone for the online class. They sell 500 raffle tickets worth Php50 each. The sole lucky winner of this raffle will receive a cash of Php1000. What is your expected net gain if you buy one raffle ticket?
1
Expert's answer
2021-07-09T07:02:09-0400
"E(x)=1000\times\dfrac{1}{500}+(-50)\times\dfrac{500-1}{500}"

"=-47.9"